A father has 50 kg and son has 30 kg. what is the ratio of their inertia ?

Grade:12th pass

1 Answers

Manas Shukla
7 years ago
Its same as ratio of their masses . Since inertia is directly proportional to mass.
So mass of father / mass of son = 50/30 = 5/3